OpenStudy (anonymous):

jimmy and kayla are selling flower bulbs for a school fundraiser. Customers can buy bags of windflower bulbs and packages of crocus bulbs. Jimmy sold 3 bags of windflower bulbs and 5 packs of crocus bulbs for a total of 104$. Kayla sold 1 bag of winflower and 5 packs of crocus bulbs for 88$ find the cost of one bag of windflower and one pack of crocus bulb

OpenStudy (anonymous):

windflower blubs = w crocus bulbs = c 3w+5c = 104 w+5c = 88 w = 88-5c 3(88-5c)+5c = 104 264-15c+5c = 104 -10c = -160 c = $16 w = 88-5(16) w = $8 price for one bag of each: w+c = price 8+16 = price $24 = price

OpenStudy (anonymous):

Thank you @speedygonzales